OpenStudy (anonymous):

whoever knows how to do these please help, preferably someone knew that can show me all the work to solve these. 4x+y=18 x=y-3

OpenStudy (kymber):

What're you supposed to do?

OpenStudy (anonymous):

solve by using the substitution method

OpenStudy (asylum15):

Put x=y-3 into 4x+y=18 So 4(y-3) + y =18 Multiply out and I'm guessing you'll get a quadratic equation which you can solve for 2 x values.

Parth (parthkohli):

It's already given that x = y - 3 So, 4x + y will be 4(y - 3) + 3

OpenStudy (asylum15):

Nevermind my reply, you'll get 4y - 12 + y = 18 5y = 18 + 12 5y = 30 y = 6